# MedViT V2: Medical Image Classification with KAN-Integrated Transformers and Dilated Neighborhood Attention

### v2

@@ Abstract
Convolutional networks, Transformers, hybrid models, and Mamba-based architectures have demonstrated strong performance across various medical image classification tasks. However, these methods were primarily designed to classify clean images using labeled data. In contrast, real-world clinical data often involve image corruptions that are unique to multi-center studies and stem from variations in imaging equipment across manufacturers. In this paper, we introduce the Medical Vision Transformer (MedViTV2), a novel architecture incorporating Kolmogorov-Arnold Network (KAN) layers into the Transformer architecture for the first time, aiming for generalized medical image classification. We have developed an efficient KAN block to reduce computational load while enhancing the accuracy of the original MedViT. Additionally, to counteract the fragility of our MedViT when scaled up, we propose an enhanced Dilated Neighborhood Attention (DiNA), an adaptation of the efficient fused dot-product attention kernel capable of capturing global context and expanding receptive fields to scale the model effectively and addressing feature collapse issues. Moreover, a hierarchical hybrid strategy is introduced to stack our Local Feature Perception and Global Feature Perception blocks in an efficient manner, which balances local and global feature perceptions to boost performance. Extensive experiments on 17 medical image classification datasets and 12 corrupted medical image datasets demonstrate that MedViTV2 achieved state-of-the-art results in 27 out of 29 experiments with reduced computational complexity. MedViTV2 is 44% more computationally efficient than the previous version and significantly enhances accuracy, achieving improvements of 4.6% on MedMNIST, 5.8% on NonMNIST, and 13.4% on the MedMNIST-C benchmark.
@@
Computer-aided diagnosis (CAD) systems have attracted significant research interest in medical image analysis, aiming to assist clinicians in making diagnostic decisions. These systems are applied to various modalities, including X-ray radiography [1], computed tomography (CT) [2], magnetic resonance imaging (MRI) [3], ultrasound [4], and digital pathology [5]. The success of deep learning in this domain is partly attributed to the increasing availability of large-scale datasets. Large datasets with reliable labels are ideal for training deep neural networks. Although, collecting labeled medical images remains challenging due to data privacy concerns and the time-consuming nature of expert annotations.
@@
Our code is available at https://github.com/Omid-Nejati/MedViTV2.git
@@ 2 Related Works
More recently, Vision Transformers (ViTs) have emerged as a powerful alternative to conventional CNNs, achieving remarkable success in various image classification tasks [32, 33, 34, 35]. ViTs offer several advantages, including the ability to model long-range dependencies, adapt to diverse inputs, and generate attention maps that highlight critical regions within an image [36]. These features have sparked significant interest in leveraging Transformer-based models for medical image classification, where precise classification is increasingly essential to support timely clinical decision-making, particularly for difficult cases. Early ViT models typically rely on large-scale datasets and relatively simple configurations [37]. However, recent advancements have integrated inductive biases related to visual perception into ViT architectures [38, 39, 40, 41]. This evolution has made ViTs more adaptable and effective in classification, registration, and segmentation. By treating images as sequences of patches without incorporating 2D inductive biases, ViTs are particularly suitable for multimodal applications [42, 43]. In particular, the growth of datasets and innovations in model architectures have driven ViT-based foundation models with unprecedented capabilities, enabling flexible applications in medical imaging [44, 45, 46]. For example, researchers have introduced FastGlioma, a tool for detecting tumor infiltration during surgery [47], while RETFound learns generalizable representations from unlabelled retinal images, enabling label-efficient model adaptation across various applications [48]. Additional applications include leveraging tumor registry data and demographic information to predict overall survival rates [49].
@@ 2.2 Kolmogorov–Arnold Networks
KANs have inspired numerous studies that demonstrate their effectiveness across various domains, including keyword spotting, complex optimization problems, survival analysis, time series forecasting, quantum computing, and vision tasks [50, 51, 52]. Furthermore, many advanced KAN models leverage well-established mathematical functions, particularly those adept at handling curves, such as B-splines [53], which combine B-splines and Radial Basis Functions (RBF) to fit input data during training. FastKAN [54] approximates third-order B-splines in KANs using Gaussian RBF, while DeepOKAN [abueidda2024deepokan] directly employs Gaussian RBF instead of B-splines. Other approaches, such as FourierKAN-GCF [xu2024fourierkan], wavelet-based KANs [seydi2024unveiling], and polynomial-function-based KANs [58], explore the suitability of various basis functions in KAN models for classification and other tasks.
@@ 3.5 MedViTV2 Architectures
To ensure a fair comparison with existing SOTA networks in medical domain, we introduce four representative variants, MedViT-V2-T/S/B/L. The architectural specifications for these variants are detailed in Table 3, where represents the stride of each stage and denotes the output channel. The spatial reduction ratio in GFP is set to [8, 4, 2, 1], while the channel shrink ratio is consistently fixed at 0.75. The expansion ratios are set to 2 for the KAN layer and and 3 for the feature expansion in LFFN. The head dimension in MHCA and E-MHSA is fixed at 32. Both LFP and GFP employ ReLU as the activation function and BatchNorm as the normalization layer.
@@ 4.5 Performance on NonMNIST
Table 7 shows the performance of MedViTV2-small and SOTA models on Kvasir and CPN X-ray. On CPN X-ray, MedViTV2-small with the lowest FLOPs achieves the best performance metrics among all models. Compared with recent models, the OA of MedViTV2-small increases by 0.9% (MedMamba-S), 1.4% (VMamba-S), 2.8% (Swin-S), and 2.6% (ConvNext-S), respectively, while maintaining a competitive parameter size. Similarly, the performance of MedViTV2-small is remarkable on Kvasir. In terms of OA, MedViTV2-small outperforms all reference models by achieving a significant improvement of 3.5% (MedMamba-T), 5.5% (VMamba-S), and 8.0% (ConvNext-S). Regarding AUC, MedViTV2-small achieves the highest result, outperforming all competitors, including the previously top-ranked Deit-small by 0.7%.
@@ 5.1 Impact of Components
In stark contrast to the fixed receptive field growth exhibited by NAT and Swin, DiNA’s receptive field demonstrates remarkable adaptability and is contingent on the chosen dilation. Its receptive field can span from a linear progression, similar to NAT’s (when all dilation values are set to 1), to an exponential growth of (when dilation increases progressively). This inherent flexibility in dilation is a primary driver of DiNA’s superior performance. Regardless of the dilation strategy, the first layer of DiNA invariably yields a receptive field of size . When sufficiently large dilation values are utilized, each DiNA layer contributes a receptive field of size per dimension, culminating in an effective receptive field of . This allows for DiNA and NA combinations, when paired with optimal dilation values, to achieve an impressive exponential receptive field expansion up to . This observation is consistent with established knowledge regarding dilated convolutions, which are known to exhibit exponential receptive field growth when using exponentially increasing dilation factors [21]. In our experiments, DiNA demonstrates notable improvements, surpassing Swin and NAT by 4.8% and 1.6%, respectively, on TissueMNIST. Furthermore, DiNA significantly excels on the corruption benchmark, outperforming Swin and NAT by 14.8% and 5.1%, respectively, on TissueMNIST-C.

### v3

# Medical Image Classification with KAN-Integrated Transformers and Dilated Neighborhood Attention

@@ Abstract
Convolutional networks, Transformers, hybrid models, and Mamba-based architectures have demonstrated strong performance across various medical image classification tasks. However, these methods were primarily designed to classify clean images using labeled data. In contrast, real-world clinical data often involve image corruptions that are unique to multi-center studies and stem from variations in imaging equipment across manufacturers. In this paper, we introduce the Medical Vision Transformer (MedViTV2), a novel architecture incorporating Kolmogorov-Arnold Network (KAN) layers into the Transformer architecture for the first time, aiming for generalized medical image classification. We have developed an efficient KAN block to reduce computational load while enhancing the accuracy of the original MedViT. Additionally, to counteract the fragility of our MedViT when scaled up, we propose an enhanced Dilated Neighborhood Attention (DiNA), an adaptation of the efficient fused dot-product attention kernel capable of capturing global context and expanding receptive fields to scale the model effectively and addressing feature collapse issues. Moreover, a hierarchical hybrid strategy is introduced to stack our Local Feature Perception and Global Feature Perception blocks in an efficient manner, which balances local and global feature perceptions to boost performance. Extensive experiments on 17 medical image classification datasets and 12 corrupted medical image datasets demonstrate that MedViTV2 achieved state-of-the-art results in 27 out of 29 experiments with reduced computational complexity. MedViTV2 is 44% more computationally efficient than the previous version and significantly enhances accuracy, achieving improvements of 4.6% on MedMNIST, 5.8% on NonMNIST, and 13.4% on the MedMNIST-C benchmark. Code is available at https://github.com/Omid-Nejati/MedViTV2.git
@@
Computer-aided diagnosis (CAD) systems have attracted significant research interest in medical image analysis, aiming to assist clinicians in making diagnostic decisions. These systems are applied to various modalities, including X-ray radiography [1], computed tomography (CT) [2], magnetic resonance imaging (MRI) [3], ultrasound [4], and digital pathology [5]. The success of deep learning in this domain is partly attributed to the increasing availability of large-scale datasets. Large datasets with reliable labels are ideal for training deep neural networks. However, the collection of labeled medical images remains a significant challenge due to data privacy concerns and the labor-intensive nature of expert annotation.
@@ 2 Related Works
More recently, Vision Transformers (ViTs) have emerged as a powerful alternative to conventional CNNs, achieving remarkable success in various image classification tasks [32, 33, 34, 35]. ViTs offer several advantages, including the ability to model long-range dependencies, adapt to diverse inputs, and generate attention maps that highlight critical regions within an image [36]. These features have sparked significant interest in leveraging Transformer-based models for medical image classification, where precise classification is increasingly essential to support timely clinical decision-making, particularly for difficult cases. Early ViT models typically rely on large-scale datasets and relatively simple configurations [37]. However, recent advancements have integrated inductive biases related to visual perception into ViT architectures [38, 39, 40, 41]. This evolution has made ViTs more adaptable and effective in classification, registration, and segmentation. By treating images as sequences of patches without incorporating 2D inductive biases, ViTs are particularly suitable for multimodal applications [42, 43]. In particular, the growth of datasets and innovations in model architectures [44, 45, 46] have driven ViT-based foundation models with unprecedented capabilities, enabling flexible applications in medical imaging [47, 48, 49]. For example, researchers have introduced FastGlioma, a tool for detecting tumor infiltration during surgery [50], while RETFound learns generalizable representations from unlabelled retinal images, enabling label-efficient model adaptation across various applications [51]. Additional applications include leveraging tumor registry data and demographic information to predict overall survival rates [52].
@@ 2.2 Kolmogorov–Arnold Networks
KANs have inspired numerous studies that demonstrate their effectiveness across various domains, including keyword spotting, complex optimization problems, survival analysis, time series forecasting, quantum computing, and vision tasks [53, 54, 55]. Furthermore, many advanced KAN models leverage well-established mathematical functions, particularly those adept at handling curves, such as B-splines [56], which combine B-splines and Radial Basis Functions (RBF) to fit input data during training. FastKAN [57] approximates third-order B-splines in KANs using Gaussian RBF, while DeepOKAN [58] directly employs Gaussian RBF instead of B-splines. Other approaches, such as FourierKAN-GCF [59], wavelet-based KANs [60], and polynomial-function-based KANs [61], explore the suitability of various basis functions in KAN models for classification and other tasks.
@@ 3.5 MedViTV2 Architectures
The overall architecture of MedViTV2 is already shown in Figure 2, while the detailed procedure is presented in Algorithm 1. To ensure a fair comparison with existing SOTA networks in medical domain, we introduce four representative variants, MedViT-V2-T/S/B/L. The architectural specifications for these variants are detailed in Table 3, where represents the stride of each stage and denotes the output channel. The spatial reduction ratio in GFP is set to [8, 4, 2, 1], while the channel shrink ratio is consistently fixed at 0.75. The expansion ratios are set to 2 for the KAN layer and and 3 for the feature expansion in LFFN. The head dimension in MHCA and E-MHSA is fixed at 32. Both LFP and GFP employ ReLU as the activation function and BatchNorm as the normalization layer.
@@ 4.5 Performance on NonMNIST
Table 7 shows the performance of MedViTV2-small and SOTA models on Kvasir and CPN X-ray. On CPN X-ray, MedViTV2-small achieves the best performance metrics among all models. Compared with recent models, the OA of MedViTV2-small increases by 0.9% (MedMamba-S), 1.4% (VMamba-S), 2.8% (Swin-S), and 2.6% (ConvNext-S), respectively, while maintaining a competitive parameter size. Similarly, the performance of MedViTV2-small is remarkable on Kvasir. In terms of OA, MedViTV2-small outperforms all reference models by achieving a significant improvement of 3.5% (MedMamba-S), 5.5% (VMamba-S), and 8.0% (ConvNext-S). Regarding AUC, MedViTV2-small achieves the highest result, outperforming all competitors, including the previously top-ranked Deit-small by 0.7%.
@@ 5.1 Impact of Components
In stark contrast to the fixed receptive field growth exhibited by NAT and Swin, DiNA’s receptive field demonstrates remarkable adaptability and is contingent on the chosen dilation. Its receptive field can span from a linear progression, similar to NAT’s (when all dilation values are set to 1), to an exponential growth of (when dilation increases progressively). This inherent flexibility in dilation is a primary driver of DiNA’s superior performance. Regardless of the dilation strategy, the first layer of DiNA invariably yields a receptive field of size . When sufficiently large dilation values are utilized, each DiNA layer contributes a receptive field of size per dimension, culminating in an effective receptive field of . This allows for DiNA and NA combinations, when paired with optimal dilation values, to achieve an impressive exponential receptive field expansion up to . This observation is consistent with established knowledge regarding dilated convolutions, which are known to exhibit exponential receptive field growth when using exponentially increasing dilation factors [21]. In our experiments, DiNA demonstrates notable improvements, surpassing Swin and NAT by 4.8% and 1.6%, respectively, on TissueMNIST. Furthermore, DiNA significantly excels on the corruption benchmark, outperforming Swin and NAT by 14.8% and 5.4%, respectively, on TissueMNIST-C.